OpenStudy (anonymous):

4.) Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

What is the slope of the line perpendicular to y = 4x + 7?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

That's the slope of the given line. What's the slope of the line perpendicular?

OpenStudy (amistre64):

id start out with a parallel line, then just trick out the slope

OpenStudy (anonymous):

Find that slope then do (y - -3 / x - 2) = m

OpenStudy (anonymous):

how do i find that slope?

OpenStudy (anonymous):

The rule is the perpendicular is the negative inverse First slope = 4 Second slope = - 1/4

OpenStudy (amistre64):

the material that is in the chapter before this questions gives you properties regarding perp slopes i beileve :/

OpenStudy (anonymous):

Okay.

OpenStudy (amistre64):

" (y - -3 / x - 2) = m " this seems a bit odd to me

OpenStudy (amistre64):

i see it now, its just the point slope form reaaranged

OpenStudy (anonymous):

i dont get what the slope is though :(

OpenStudy (amistre64):

when given a slope: a/b the perp slope is simply the flip negation of it: -b/a

OpenStudy (amistre64):

you say correctly that the slope given is 4, lets say 4/1 can you show me the perp slope from that?

OpenStudy (anonymous):

whats a perp slope??

OpenStudy (amistre64):

in the context of the question at hand, i would say that perp slope is shorthand for perpendicular slope

OpenStudy (amistre64):

mathikers hate to write :)

OpenStudy (amistre64):

here is a fairly simple method step 1, write out the given equation y = 4x + 7 ignore the constant, we are going to have to resolve for it y = 4x + b perp the slope y = (-1/4)x + b this is almost done now. plug in the values of x and y from the given point (2,-3) to solve for b -3 = (-1/4)2 + b -3 = -1/2 + b -3 + 1/2 = b (-6 + 1)/2 = b -5/2 = b plug in the value of b back into the equation y = (-1/4)x - 5/2 and voila

OpenStudy (amistre64):

thats not all step 1.... i just forgot to write in "step ... " for the others lol

OpenStudy (anonymous):

it makes so much sense now! and ok lol. thank you very!!!!!!!!!!!!

OpenStudy (amistre64):

the term "standard form" youll have to google; it might be what we did, or it might mean: ax + by = c format, i cant recall

OpenStudy (anonymous):

your correct!

OpenStudy (amistre64):

yay!! staving off althiemers for one more day :)

OpenStudy (anonymous):

lol thanks again (: